OpenStudy (anonymous):

If f(x) = x + 3, find f(-8). -11 -5 5 11

OpenStudy (lgbasallote):

substitute -8 into x f(-8) = - 8 + 3

OpenStudy (anonymous):

Didn't you get the answer @Ashley16 ??

OpenStudy (anonymous):

no..

OpenStudy (lgbasallote):

-8 + 3.....

OpenStudy (lgbasallote):

combine

OpenStudy (anonymous):

What you get when you subtract 3 from 8 @Ashley16 ??

OpenStudy (anonymous):

\[\huge \color{green}{-8 + 3 = -(8-3)}\] \[\huge \color{red}{8 - 3 = ??}\]

OpenStudy (anonymous):

5 i know, i know.

OpenStudy (anonymous):

But the answer is not 5 you know that??

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome dear..